OpenStudy (anonymous):

A manufacturer of bird feeders cuts spacers from a tube that is 7 5/7 inches long. How many spacers can be cut from the tube if each spacer must be 6/7 inch? estimated number of spacers? exact number of spacers?

OpenStudy (anonymous):

(7 5/7) / (6/7)

OpenStudy (anonymous):

25/2 ?

OpenStudy (anonymous):

which would be 12.5

OpenStudy (anonymous):

estimate 12 and actual 13?

ganeshie8 (ganeshie8):

in the question, is it \(7\frac{5}{7}\) or \(\frac{75}{7}\) ?

OpenStudy (anonymous):

the first one

ganeshie8 (ganeshie8):

then you need to convert the mixed fraction

OpenStudy (anonymous):

so was i right with my answer earlier?

ganeshie8 (ganeshie8):

here is an insane way of doing it : 7 5/7 7 + 5/7 7*7/7 + 5/7 [7(6+1)/7) + 5/7] 7(6/7) + 7/7 + 5/7 7(6/7) + 6/7 + 1/7 + 5/7 7(6/7) + 6/7 + 6/7 9(6/7)

ganeshie8 (ganeshie8):

so there are 9 objects of length 6/7 !!

OpenStudy (anonymous):

ok I see now by your breaksown

ganeshie8 (ganeshie8):

great !! but the easy way is to simply divide the 7 5/7 with 6/7 as below 7 5/7 = 54/7 to findout spacers of length 6/7 : (54/7)/(6/7) = 54/6 = 9 both methods give the same answer